### v5

# Revisit the Arimoto-Blahut algorithm: New Analysis with Approximation

Abstract
By the seminal paper of Claude Shannon [Shannon48], the computation of the capacity of a discrete memoryless channel has been considered as one of the most important and fundamental problems in Information Theory. Nearly 50 years ago, Arimoto and Blahut independently proposed identical algorithms to solve this problem in their seminal papers [Arimoto1972AnAF, Blahut1972ComputationOC]. The Arimoto-Blahut algorithm was proven to converge to the capacity of the channel as $t\to\infty$ with the convergence rate upper bounded by $O\left(\log(m)/t\right)$, where $m$ is the size of the input distribution, and being inverse exponential when there is a unique solution in the interior of the input probability simplex [Arimoto1972AnAF]. Recently it was proved, in [Nakagawa2020AnalysisOT], that the convergence rate is at worst inverse linear $O(1/t)$ in some specific cases.
In this paper, we revisit this fundamental algorithm looking at the rate of convergence to the capacity and the time complexity, given $m,n$, where $n$ is size of the output of the channel, focusing on the approximation of the capacity. We prove that the rate of convergence to an $\varepsilon$-optimal solution, for any sufficiently small constant $\varepsilon>0$, is inverse exponential $O\left(\log(m)/c^{t}\right)$, for a constant $c>1$ and $O\left(\log\left(\log(m)/\varepsilon\right)\right)$ at most iterations, implying $O\left(mn\log\left(\log(m)/\varepsilon\right)\right)$ total complexity of the algorithm.
1 Introduction
In Information theory [Cover2006], the Arimoto-Blahut algorithm stands as a cornerstone method for efficiently computing the capacity of discrete memoryless channels and solving related optimization problems. Originally introduced by Suguru Arimoto and Richard Blahut in the 1970s [Arimoto1972AnAF, Blahut1972ComputationOC], and it has been celebrated for its convergence properties and practical applicability in diverse communication systems. In particular, a series of input distributions are computed by a recurrence formula produced by an alternating optimization approach, converging to the input probability distribution that achieves the channel capacity. In the paper of Arimoto (see also [SutterSEL15, Jurgensen1984ANO]), this algorithm was proved to have an upper bound of the rate of convergence equals to $O\Big{(}\log(m)/{t}\Big{)}$, which implies an $O(mn\log(m)/{\varepsilon})$ complexity to achieve an $\varepsilon$-optimal solution, where $m$ is the size of the input and $n$ is the size of the output. However, it is an open question if a better upper bound holds for the rate of convergence to the computation of the capacity.
Recently, this algorithm has been scrutinized heavily in [Nakagawa2020AnalysisOT], where the authors perform Taylor expansions and eigenvalues’ analysis of a Jacobian matrix to determine the conditions for the convergence rates of the algorithm for an optimal solution, showing that the algorithm in some cases converges with exponential, but in other cases with inverse linear convergence rate under some specific conditions. Furthermore, in another recent work by [SutterSEL15], utilizing the dual formulation for computing capacity, it has been demonstrated an a priori error bound of $O\left({n\sqrt{\log(m)}}/{t}\right)$. This results in overall computational complexity of $O\left({mn^{2}\sqrt{\log(m)}}/{\varepsilon}\right)$ for finding an additive $\varepsilon$-solution.
In this paper, we take a different point of view, observing that in case the algorithm is initialized in the interior of the probability space, the sequence of the probability distributions that are generated by the algorithm will converge with an inverse exponential rate to an approximate optimal solution from the interior for any constant approximation parameter $\varepsilon$. The intuition behind this is that we split the region of convergence into two regions depending on the constant approximation $\varepsilon$ of the capacity: a region R around the convex set of the optimal solutions containing all $\varepsilon$-optimal input distributions and the complement of this. This separation helped us to prove that for any constant $\varepsilon>0$ the inverse exponential rate of convergence holds until we reach the region R, or in other words the Arimoto-Blahul algorithm is exponential fast to reach a constant approximation of the capacity, for any fixed constant approximation. Furthermore, applying tools from this approach we can prove that if the convex set of the optimal solutions is bounded by below by a constant, the same bounds hold for achieving an optimal solution.
Our contributions in a more technical perspective are summarized as follows:
  • Firstly, the sequence of approximations that is generated by the Arimoto-Blahut algorithm converges to an $\varepsilon$-optimal solution, for any constant $\varepsilon>0$ (Theorem 3).
  • Secondly, the rate of convergence to an $\varepsilon$-optimal solution is upper bounded by $O\left({\log(m)}/{c^{t}}\right)$, for a $c>1$, for any $t>0$, which implies at most $O\left(\log\left({\log(m)}/{\varepsilon}\right)\right)$ iterations (Theorem 4).
  • If the region of the optimal solutions has a strictly positive volume, then the same convergence bounds apply for achieving an optimal solution (Theorem 5).
These contributions represent a significant improvement not only over the previously established upper bounds of $O\left({\log(m)}/{t}\right)$ for the convergence rate and of $O\left(mn{\log(m)}/{\varepsilon}\right)$ for the complexity for the algorithm, but also compared to the bounds of $O\left({\sqrt{\log(m)}}/{t}\right)$ and $O\left({n^{2}m\sqrt{\log(m)}}/{\varepsilon}\right)$, respectively, of the algorithm derived in [SutterSEL15].
1.1 Further Related work
Different versions of the Arimoto-Blahut algorithm are studied in bibliography. Namely, authors in [Sayir2000IteratingTA] studied the case where a small number of the input symbols are assigned non-zero probabilities in the capacity-achieving distribution, but no theoretical results regarding the speed up and the convergence rate are provided. In [Yu2009SqueezingTA], accelerated versions of the classical Arimoto-Blahut algorithm are provided, by parametrizing the recurrence formula of classical algorithm, and their work demonstrated that the convergence is at least as fast as the Arimoto-Blahut algorithm [Arimoto1972AnAF]. Moreover, authors in [Chen2023ACB] introduced a modification to the classical Arimoto-Blahut algorithm and examined the dual problem of computing the channel capacity, known as the rate-distortion problem. They reported a convergence rate of at least $O(1/t)$ and derived that their algorithm requires at least ${2\log(n)}/{\varepsilon}$ iterations. This results in $O\left(\frac{mn\log(n)}{\varepsilon}\left(1+\log|\log(\varepsilon)|\right)\right)$ arithmetic operations to achieve an $\varepsilon$-optimal solution. Further, in [Matz2004InformationGF] authors provided a modified version of the Arimoto-Blahut algorithm utilizing natural gradient and proximal point methods, and they proved the convergence of their algorithm but their convergence rate analysis is restricted only for the worst case, which is $O(1/t)$.
2 Preliminaries
We consider a discrete memoryless channel $X\to Y$, where $X$ is the input of size $m$ and $Y$ of size $n$ is the output random variables, with $\{x_{1},x_{2},\dots,x_{m}\}$ being the possible symbols of the input and $\{y_{1},y_{2},\dots,y_{n}\}$ being the possible symbols of the output. Given a probability distribution $p$, the support of $p$ is the set of indexes with positive probability. Furthermore, we consider the channel matrix $W$ with $w_{ij}\in[0,1]$ where $w_{ij}$ is the conditional probability of the $j$ output symbol when the input symbol $i$ is transmitted, with $\sum_{j\in[m]}w_{ij}=1$ for any possible $i$. Thus, for any probability distribution $p$ of the input, the output probability distribution is $q^{T}=p^{T}\cdot W$.
Further, the Shannon entropy for a distribution $p$ is defined as $H(p)=-\sum_{i\in[n]}p_{i}\cdot\log(p_{i})$, the mutual information given the channel $W$ and an input probability $p$ is defined as
$I(p,W)=\sum_{i\in[n]}\sum_{j\in[m]}p_{i}\cdot w_{ij}\log\left(\frac{w_{ij}}{\sum_{k\in[n]}p_{k}\cdot w_{kj}}\right),$
and the channel capacity $C^{*}$ is defined as $C^{*}=\max_{p\in\Delta(X)}I(p,W)$, where $\Delta(X)$ is the simplex of the input probability distribution. An input distribution that maximizes the capacity is symbolized as $p^{*}$. In general, this optimal distribution is not unique and as was proven in [Arimoto1972AnAF], any convex combination of optimal solutions is also an optimal solution. Thus, the set of optimal solutions forms a convex subset of $\Delta(X)$.
Additionally, given a set $S$ we denote as $S^{o}$ and $\partial S$ the interior and the boundary sets of $S$, respectively, and we also define the $\delta$-ball around a $p\in S$ as $B_{S}(p,\delta):=\{q\in S\mid\|p-q\|_{1}\leq\delta\}$ (or simply as $B(p,\delta)$ when the set is obvious from the context).
The Kullback-Leibler (KL) divergence $D_{KL}(p||q)$ for two distributions $p,q\in\Delta(X)$ is defined by $D_{KL}(p||q)=\sum_{i\in[m]}p_{i}\log\Big{(}\tfrac{p_{i}}{q_{i}}\Big{)}$. Then, the difference of the KL divergences between $p^{*}$ and two consecutive probability approximations $p^{t}$ and $p^{t+1}$, in time $t$ and $t+1$, respectively, as derived by the algorithm is given by
$D_{KL}(p^{*}||p^{t+1})-D_{KL}(p^{*}||p^{t})=\sum\nolimits_{i\in[m]}p^{*}_{i}\log\left(\frac{p^{t}_{i}}{p^{t+1}_{i}}\right),\quad\text{for }t\geq 0.$ (1)
Furthermore, in [Nakagawa2020AnalysisOT] the authors classify the indexes of an optimal solution $p^{*}$, since in the maximization problem $\max_{p\in\Delta(X)}I(p,W)$, a necessary and sufficient condition for an optimal input distribution $p^{*}=(p^{*}_{1},\ldots,p^{*}_{m})\in\Delta(X)$ to achieve the capacity of the channel is the Karush-Kuhn-Tucker conditions, see [Jelinek1968ProbabilisticIT], thus we have that:
$D(W^{i}||q^{*})\left\{\begin{array}[]{ll}=C^{*},&p^{*}_{i}>0,\\ \leq C^{*},&p^{*}_{i}=0,\\ \end{array}\right.$
where $(q^{*})^{T}=(p^{*})^{T}\cdot W$ and $W^{i}$ is the $i$-th row of the matrix $W$. In other words, if the index $i$ is in the support of the optimal solution $p^{*}$, then $D(W^{i}||q^{*})=C^{*}$, but there is the case for an index $i$ that is not in the support of $p^{*}$ to achieve capacity too. Thus, as in [Nakagawa2020AnalysisOT] we split the indexes $[m]$ into:
  • Type I: For $i\in[m]$ such that $D(W^{i}||q^{*})=C^{*}$ and $p^{*}_{i}>0$,
  • Type II: For $i\in[m]$ such that $D(W^{i}||q^{*})=C^{*}$ and $p^{*}_{i}=0$,
  • Type III: For $i\in[m]$ such that $D(W^{i}||q^{*})<C^{*}$ and $p^{*}_{i}=0$.
For notational convenience, we denote as $m_{I}$, $m_{II}$, and $m_{III}$ the set of indexes of Type I, Type II, and Type III, respectively, in an optimal solution $p^{*}$.
2.1 Description of the Arimoto-Blahut algorithm
At the original Arimoto-Blahut algorithm, the time is split into discrete slots $t\in\mathbb{N}$. We consider $C(t+1,t)$ as the approximate channel capacity at the iteration $t+1$ given the probability distribution $p^{t}$ from the previous iteration. Formally, this can be written as follows
$\begin{split}C(t+1,t)=-\sum_{i\in[m]}p^{t+1}_{i}\cdot\log(p^{t+1}_{i})+\sum_{i\in[m]}\sum_{j\in[n]}p^{t+1}_{i}\cdot w_{ij}\cdot\log\left(\frac{p^{t}_{i}\cdot w_{ij}}{\sum_{k\in[n]}p^{t}_{k}\cdot w_{kj}}\right).\end{split}$ (2)
Note that at time $t+1$, in order to compute the $C(t+1,t)$, the new probability distribution $p^{t+1}$, given the previous distribution $p^{t}$, is updated by the following recurrence formula, see [Arimoto1972AnAF, Nakagawa2020AnalysisOT],
$\begin{split}p^{t+1}_{i}=\frac{p^{t}_{i}\cdot e^{\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)}}{{\sum}\nolimits_{\ell\in[m]}p^{t}_{\ell}\cdot e^{\sum_{j\in[n]}w_{\ell j}\cdot\log\left(\frac{w_{\ell j}}{\sum_{k\in[n]}p^{t}_{k}\cdot w_{kj}}\right)}},\quad\text{for any }i\in[m],\end{split}$ (3)
assuming that $p^{0}$ is a full support uniform distribution, in other words $p^{0}_{i}=1/m$ for any $i$. Noteworthy, if $p^{0}_{i}>0$ then $p^{t}_{i}>0$, for any $t$, that is if we start from a point in the interior of $\Delta(X)$ the convergence path will remain in the interior.
2.2 Approximation of the capacity
Let the function $f:\Delta(X)\to[0,C^{*}]$ be defined by $f(p^{t})=C^{*}-C(t+1,t)$ representing the distance/error from the capacity at time $t+1$, such that the problem of the computation of the capacity of the channel becomes a minimization problem, where the minimum value is zero for any optimal distribution $p^{*}$. Now, we have the following approximation of the function $f$.
Definition 1 (Approximate optimum solution).
A distribution $p$ is $\varepsilon$-optimum solution, if and only if,
$f(p)\leq\varepsilon,\quad\text{ for any }\varepsilon\geq 0.$
It is easy to see that when $\varepsilon=0$ we have an (exact) optimum solution. This leads us to the following straightforward result.
Lemma 1.
For any optimum solution $p^{*}$ there is a set of solutions $S\in\Delta^{o}(X)$, s.t. for any $p\in S$ it holds $f(p)\leq\varepsilon$, for any sufficiently small $\varepsilon>0$ and $||p-p^{*}||^{2}_{2}\leq\delta(\varepsilon)$, for a sufficiently small $\delta(\varepsilon)>0$.
Proof.
The proof is straightforward by the continuity of the function $f$. ∎
We denote the interior of the set of $\varepsilon$-optimal solutions as $B^{o}(p^{*},\delta(\varepsilon))$, for any $p^{*}$ optimum solution. The previous Lemma implies that for any optimal solution $p^{*}$, even for solutions on the boundary $\partial\Delta(X)$, that is with $m_{II}\neq\emptyset$, there is a set of approximate solutions $B^{o}(p^{*},\delta(\varepsilon))$ with $m_{II}=\emptyset$ and $f(p)\leq\epsilon$ for any $p\in B^{o}(p^{*},\delta(\varepsilon))$. Moreover, in this paper, we define two types of convergence rates for the function $f$:
Definition 2 (Inverse linear rate).
The function $f$ has inverse linear rate of convergence, as $f(p^{t})\to\varepsilon$, for any $\varepsilon\in[0,C^{*}]$, if and only if,
$f(p^{t})=O(1/t),$
for any $t\in\mathbb{Z^{+}}$, provided that $f(p^{t})>\varepsilon$.
Definition 3 (Inverse exponential rate).
The function $f$ has inverse exponential rate of convergence, as $f(p^{t})\to\varepsilon$, for any $\varepsilon\in[0,C^{*}]$, if and only if,
$f(p^{t})=O(1/c^{t}),$
for any $t\in\mathbb{Z^{+}}$, with $c>1$, provided that $f(p^{t})>\varepsilon$.
Given these definitions, the previous main results of the bibliography on the rate convergence of the Arimoto-Blahut algorithm are summarized as following:
Corollary 1.
[Corollary 1 in [Arimoto1972AnAF]] The Arimoto-Blahut algorithm has rate of convergence for $f(p^{t})$ at most $O\Big{(}{\log(m)}/{t}\Big{)}$ to converge to an optimal solution.
This corollary implies that the Arimoto-Blahut algorithm requires at most $O\left({\log(m)}/{\varepsilon}\right)$ iterations to find an $\varepsilon$-optimal solution [SutterSEL15, Jurgensen1984ANO].
Theorem 1.
[Theorem 5 in [Nakagawa2020AnalysisOT] and Theorem 3 in [Arimoto1972AnAF]] Consider a unique optimal solution with $m_{II}=\emptyset$, then the Arimoto-Blahut algorithm has inverse exponential rate of convergence [^1] For the capacity or the distance/norm of the current distribution from the optimum. We can consider these two metrics as ”equivalent”. to achieve this optimal solution.
Theorem 2.
[Theorem 9 in [Nakagawa2020AnalysisOT]] If $m_{II}\neq\emptyset$ in the unique optimal solution, then there is an initial distribution $p^{0}$ with $p^{0}_{i}>0$, for any $i\in m_{II}$, s.t. the Arimoto-Blahut algorithm has inverse linear rate of convergence $O(1/t)$ to converge to this optimal solution as $t\to\infty$.
The previous Theorem gives an upper bound to the rate of convergence to an optimal solution. To improve this bound, we now consider an approximation of the optimal solutions.
3 Approximation analysis of the Arimoto-Blahut algorithm
We start out the analysis by proving that the sequence of the distributions generated by the Arimoto-Blahut algorithm (3) converges to an $\varepsilon$-optimal solution $p\in B(p^{*},\delta(\varepsilon))$ for some optimal solution $p^{*}$, as far as $f>\varepsilon$, for any constant $\varepsilon>0$. Formally, this is stated as:
Theorem 3.
Consider the sequence of distributions $\{p^{t}\}_{t\in\mathbb{N^{+}}}$ generated by equations (3), starting from a full support uniform initial distribution $p^{0}$. Then, $\{p^{t}\}_{t\in\mathbb{N^{+}}}$ converges to an $\varepsilon$-optimal solution, for any constant $\varepsilon>0$, as far as $f(p^{t})>\varepsilon$.
Proof.
The proof of this Theorem closely follows the approach used in equations (30) - (32) in [Arimoto1972AnAF], but we present it again for the sake of completeness of the analysis. Since the set of the optimal solutions $p^{*}$ is convex, we can pick an arbitrary optimal solution $p^{*}$, in other words, no matter what optimal solution we pick we can prove that the algorithm always converges close to the convex set of the optimum solutions that is an $\varepsilon$-optimum solution as far as $f(p^{t})>\varepsilon$.
Since, we consider that the initial distribution $p^{0}$ is a full support uniform distribution, the Arimoto-Blahut algorithm is guaranteed to converge to a solution $p\in B(\bar{p},\delta(\varepsilon))$, for an optimum solution $\bar{p}$, through a path [^2] This happens since any index in any approximation that is generated from the algorithm has a strictly positive probability.in the $\Delta^{o}(X)$, in other words any $p^{t}\in\Delta^{o}(X)$. To prove this, we calculate the difference of the KL divergences between the optimal distribution $p^{*}$ and two consecutive approximations, $p^{t}$ and $p^{t+1}$ and the optimum solution $p^{*}$, proving that it is positive, that is the KL distance between the approximation and the $p^{*}$ decreases with increasing $t$. Hence, the sequence approaches the set of the optimal solutions as long as $f(p^{t})>\varepsilon$. Thus, we have that
$\begin{split}D_{KL}(p^{*}||p^{t+1})-D_{KL}(p^{*}||p^{t})&=\sum\nolimits_{i\in[m]}p^{*}_{i}\log\left(\frac{p^{t}_{i}}{p^{t+1}_{i}}\right)\\ &=\sum_{i\in[m]}p^{*}_{i}\cdot\log\left(\frac{p^{t}_{i}\cdot{\sum}\nolimits_{\ell\in[m]}e^{\sum_{j\in[n]}w_{\ell j}\cdot\log\left(\frac{p^{t}_{\ell}\cdot w_{\ell j}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)}}{p^{t}_{i}\cdot e^{\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)}}\right)\\ &=\log\left({\sum}_{\ell\in[m]}e^{\sum_{j\in[n]}w_{\ell j}\cdot\log\left(\frac{p^{t}_{\ell}\cdot w_{\ell j}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)}\right)\\ &\hskip 50.00008pt-{\sum}_{i\in[m]}p^{*}_{i}\cdot\log\left(e^{\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)}\right).\\ \end{split}$
Setting $r^{t}_{i}=e^{\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{p^{t}_{i}\cdot w_{ij}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)}$ and using the properties of the logarithms we have
$D_{KL}(p^{*}||p^{t+1})-D_{KL}(p^{*}||p^{t})=\log\left(\sum_{i\in[m]}r^{t}_{i}\right)-\sum_{i\in[m]}p^{*}_{i}\cdot\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right).$
Using that $C(t+1,t)=\log\Big{(}\sum\nolimits_{i\in[m]}r^{t}_{i}\Big{)}$ (see Corollary 1 and equation (28) in [Arimoto1972AnAF]) and rearranging terms, the KL divergences difference becomes
$\begin{split}D_{KL}(p^{*}||p^{t+1})&-D_{KL}(p^{*}||p^{t})=C(t+1,t)-\sum_{i\in[m]}p^{*}_{i}\cdot\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)\\ &=C(t+1,t)-\sum_{i\in[m]}p^{*}_{i}\cdot\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}\cdot\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}{\left(\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}\right)\cdot\left(\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}\right)}\right)\\ &=C(t+1,t)-\sum_{i\in[m]}p^{*}_{i}\cdot\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}\right)\\ &\hskip 130.0002pt-\sum_{i\in[m]}p^{*}_{i}\cdot\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right).\\ \end{split}$
Now, since $C^{*}=\sum_{i\in[m]}p^{*}_{i}\cdot\sum_{j\in[n]}w_{ij}\cdot\log\left(\frac{w_{ij}}{\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}\right)$ we have that
$D_{KL}(p^{*}||p^{t+1})-D_{KL}(p^{*}||p^{t})=C(t+1,t)-C^{*}-\sum_{i\in[m]}\sum_{j\in[n]}p^{*}_{i}\cdot w_{ij}\cdot\log\left(\frac{\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right).\$
The output of the communication channel is $q^{t}_{j}=\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}$, hence
$\sum_{i\in[m]}\sum_{j\in[n]}p^{*}_{i}\cdot w_{ij}\cdot\log\left(\frac{\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)=\sum_{j\in[n]}\sum_{i\in[m]}p^{*}_{i}\cdot w_{ij}\cdot\log\left(\frac{\sum_{k\in[m]}p^{*}_{k}\cdot w_{kj}}{\sum_{k\in[m]}p^{t}_{k}\cdot w_{kj}}\right)=D_{KL}(q^{*}||q^{t}).$
Therefore,
$D_{KL}(p^{*}||p^{t+1})-D_{KL}(p^{*}||p^{t})=C(t+1,t)-C^{*}-D_{KL}(q^{*}||q^{t})=-f(p^{t})-D_{KL}(q^{*}||q^{t}).$ (4)
Since, the KL divergence $D_{KL}(q^{*}||q^{t})$ is always non negative and $f(p^{t})>\varepsilon$, we have that $D_{KL}(p^{*}||p^{t+1})-D_{KL}(p^{*}||p^{t})\leq-f(p^{t})<-\varepsilon$, for any $p^{t}\in\Delta^{o}(X)$ such that $f(p^{t})>\varepsilon$. This implies that the KL divergence decreases by at least a constant quantity $\varepsilon$ in each iteration $t$, and the sequence $\{p^{t}\}_{t\in\mathbb{N}}$ approaches the convex set of the optimal solutions provided that we have not reached an $\varepsilon$-optimal solution and the proof is complete. ∎
We can easily see that from the proof of Theorem 3 we have that $D_{KL}(p^{*}||p^{t})-D_{KL}(p^{*}||p^{t+1})>\varepsilon$, thus $D_{KL}(p^{*}||p^{0})-D_{KL}(p^{*}||p^{t+1})>t\cdot\varepsilon$, for any $t$ s.t. $f(p^{t})>\varepsilon$. Therefore, we can take that $D_{KL}(p^{*}||p^{0})>t\cdot\varepsilon$, which implies that $t<\frac{D_{KL}(p^{*}||p^{0})}{\varepsilon}<\frac{\log(m)}{\varepsilon}$ matching the previous upper bound (see Corollary 1 in [Arimoto1972AnAF] and [SutterSEL15]) of the analysis for the Arimoto-Blahut algorithm to reach an $\varepsilon$-optimal solution.
In the following theorem, we improve the bound of the rate of convergence and the upper bound of the worst number of steps to achieve an $\varepsilon$-optimal solution, and this is the main result of our paper.
Theorem 4.
Consider the sequence of distributions $\{p^{t}\}_{t\in\mathbb{N^{+}}}$ generated by equations (3), starting from a full support uniform initial distribution $p^{0}$. As long as $f(p^{t})>\varepsilon$, for any sufficiently small constant [^3] We can also consider $\varepsilon$ as a function of the size of the problem $(m,n)$, for instance, $\varepsilon=\frac{1}{poly(m,n)}$, since in the analysis of the complexity we can consider that $(m,n)\to\infty$, but for a specific realization of the problem we treat $(m,n)$ as fixed constants even if they are arbitrarily large. $\varepsilon>0$, the rate of convergence for the function $f$ to achieve an $\varepsilon$-optimal solution is inverse exponential
$O\left(\frac{\log(m)}{c^{t}}\right),$
with $c>1$ being a constant, for any $t>0$. This rate implies at most $O\left(\log\left(\frac{\log(m)}{\varepsilon}\right)\right)$ iterations to reach an $\varepsilon$-optimal solution, and a total complexity of
$O\left(mn\log\left(\frac{\log(m)}{\varepsilon}\right)\right),$
considering that in any iteration we have a $O(mn)$ computational complexity.
Proof.
By the previous Theorem we have that $f(p^{t})=D_{KL}(p^{*}||p^{t})-D_{KL}(p^{*}||p^{t+1})-D_{KL}(q^{*}||q^{t})$. Applying now the data processing inequality, see in [Csiszr2011InformationT], for the KL divergence, we have
$D_{KL}(p^{*}||p)\geq D_{KL}({p^{*}}^{T}W||p^{T}W)=D_{KL}(q^{*}||q),$
for all $p^{*},p\in\Delta(X)$. In other words, there exist an $a_{t}\in(0,1]$ [^4] In case $a_{t}=0$ then $D_{KL}(q^{*}||q^{t})=0$, which implies that $q^{*}=q^{t}$ and we achieve the capacity. Therefore, the input distribution $p^{t}$ provides the capacity as well. such that $D_{KL}(q^{*}||q^{t})=a_{t}\cdot D_{KL}(p^{*}||p^{t})$. Since $f(p^{t})>\varepsilon$, we have
$(1-a_{t})\cdot D_{KL}(p^{*}||p^{t})-D_{KL}(p^{*}||p^{t+1})=f(p^{t})>\varepsilon,$ (5)
which implies that $a_{t}<1$ and $f(p^{t})\leq D_{KL}(p^{*}||p^{t})$. Rearranging terms in (5), then we take
$\begin{split}D_{KL}(p^{*}||p^{t+1})<(1-a_{t})\cdot D_{KL}(p^{*}||p^{t})-\varepsilon,\end{split}$ (6)
Similar we can argue that in the previous step, $t-1$, by (6) we have $D_{KL}(p^{*}||p^{t})=D_{KL}(p^{*}||p^{t-1})-\eta\cdot\varepsilon=(1-b_{t-1})\cdot D_{KL}(p^{*}||p^{t-1})$, with $\eta\geq 1$ and $b_{t-1}=\frac{\eta\cdot\varepsilon}{D_{KL}(p^{*}||p^{t-1})}\geq\frac{\varepsilon}{\log(m)}$, so we can argue that, since $\varepsilon$ is a constant and $\log m$ is a constant, that $b_{t-1}\geq\zeta_{t-1}$ where $\zeta_{t-1}$ a constant independent of $t$. Thus, from (5) we have $f(p^{t})\leq D_{KL}(p^{*}||p^{t})\leq(1-b_{t-1})\cdot D_{KL}(p^{*}||p^{t-1})\leq(1-\zeta_{t-1})\cdot D_{KL}(p^{*}||p^{t-1})$. Let now $\zeta=\min_{t<t^{*}}\{\zeta_{t}\}\in(0,1)$.
This implies that $\varepsilon<f(p^{t})\leq(1-b_{t-1})\cdot D_{KL}(p^{*}||p^{t-1})\leq(1-\zeta_{t-1})\cdot D_{KL}(p^{*}||p^{t-1})\leq(1-\zeta)\cdot D_{KL}(p^{*}||p^{t-1})$. Taking into account that $D_{KL}(p^{*}||p^{0})\leq\log(m)$, then for $t<t^{*}$ it holds $f(p^{t})<(1-\zeta)^{t}\cdot\log(m)=\frac{\log(m)}{c^{t}}$, where $c=1/(1-\zeta)>1$ be a constant. That is, the approximation error decreases exponentially to any given constant minimum value $\varepsilon>0$, and we can write, for any $t\leq t^{*}$,
$f(p^{t})=O\left(\frac{\log(m)}{c^{t}}\right).$
Furthermore, by this convergence rate, we have that the number of iterations to achieve an $\varepsilon$-optimal solution provided by the formula $\frac{\log(m)}{c^{t}}=\varepsilon$, are at most
$t=O\left(\log\left(\frac{\log(m)}{\varepsilon}\right)\right).$
Finally, since in any iteration we have at most $O(mn)$ calculations, the total complexity is of
$O\left(mn\log\left(\frac{\log(m)}{\varepsilon}\right)\right).$
Note that the previous Theorem holds for any case of optimal solutions, including the case of a unique solution with $m_{II}\neq\emptyset$, where the authors in [Nakagawa2020AnalysisOT] prove the $O(1/t)$ convergence, as $t\to\infty$, for $f(p^{t})\to 0$. According to our Theorem, in such cases, there exists a set $B^{o}(p^{*},\delta(\varepsilon))$ of $\varepsilon$-optimal solutions that the algorithm can approach via an interior path with an inverse exponential rate of convergence, see Figure LABEL:fig:thm4. However, to achieve the capacity, the algorithm, after reaching an $\varepsilon$-optimal solution exhibits an $O(1/t)$ convergence rate as $t\to\infty$ as prove in [Nakagawa2020AnalysisOT].
In other words, the key idea of our analysis is that we divide the convergence process into two phases: the initial phase to reach an $\varepsilon$-optimal solution and the subsequent phase to achieve an exact optimal solution, see Figure LABEL:fig:Convergence_of_AB. We demonstrate that the rate of convergence to an $\varepsilon$-optimal solution is always inverse exponential for any constant $\varepsilon>0$, regardless of the channel.
4 A consideration of inverse exponential rate of convergence for exact solutions
We now continue our analysis considering the following rationale. An $n$-dimensional simplex, such as $\Delta(X)$, is the convex hull of $n+1$ vertices. The interior of a simplex, $\Delta^{o}(X)$ in our case, consists of all points that can be expressed as a convex combination of its vertices, where all coefficients are strictly between 00 and $1$. Then, when a convex set $S$ intersects $\Delta^{o}(X)$, it will contain such points with strictly positive coefficients. Hence, $S^{0}$ is a non-empty set and will contain a maximum possible ball $B(x_{0},\rho)$ around a point $x_{0}$ s.t. any $p\in B(x_{0},\rho)$ is an optimal solution. We consider now the case in which this radius $\rho$ is a constant and give the following result.
Theorem 5.
Consider the simplex $\Delta(X)$ and the convex set of the optimal solutions $S$, with $B(p^{*},\rho)\in S$ be the maximum possible ball in $S$, with center $p^{*}$ and constant radius $\rho$. Then, applying the Arimoto-Blahut algorithm with a full support uniform initial distribution $p^{0}$, we will have a sequence of distributions $\{p^{t}\}_{t\in\mathbb{N^{+}}}$ that converges to an exact optimal solution in $S$ with the rate of convergence equal to $O\left({\log(m)}/{c^{t}}\right)$, for a constant $c>1$, with at most $O\left(\log\left({\log(m)}/{\varepsilon}\right)\right)$ iterations to reach an optimal solution, and with $O\left(mn\log\left({\log(m)}/{\varepsilon}\right)\right)$ total complexity, for some constant $\varepsilon>0$.
Proof.
By the assumptions $B(p^{*},\rho)$ is the greatest possible ball in $S$ with constant radius $\rho$. Then, there exists a constant $\xi$ such that $\xi=\max\{D_{KL}(q^{*}||q)\mid\text{ for any }p\in B(p^{*},\rho)\}$, with $q^{T}=p^{T}\cdot W$. Consider any sequence $\{p^{t}\}\notin S$ produced by the algorithm, then it holds that $D_{KL}(q^{*}||q^{t})>\xi>c^{*}\cdot\xi$, for a sufficiently small constant $c^{*}\in(0,1)$.
By equation (4) we have that $D_{KL}(p^{*}||p^{t})-D_{KL}(p^{*}||p^{t-1})<-D_{KL}(q^{*}||q^{t})<-c^{*}\cdot\xi$, so $D_{KL}(p^{*}||p^{t})=D_{KL}(p^{*}||p^{t-1})-\eta\cdot\xi$, for some $\eta>c^{*}$. Hence at any time $t>0$ there is a constant $\xi_{t}\in(0,1)$ such that $D_{KL}(p^{*}||p^{t})=\xi_{t}\cdot D_{KL}(p^{*}||p^{t-1})$, which implies that $\xi_{t}\cdot D_{KL}(p^{*}||p^{t-1})=D_{KL}(p^{*}||p^{t-1})-\eta\cdot\xi$. Thus, $\xi_{t}=1-\frac{\eta\cdot\xi}{D_{KL}(p^{*}||p^{t-1})}<1-\frac{c^{*}\cdot\xi}{\log(m)}$ and since $c^{*},\xi,\log(m)$ are constants independent of $t$, then $\xi_{t}$ is upper bounded by a constant less than 1. Then taking the constant $c=1/\max_{t}\{\xi_{t}\}>1$ we have that $D_{KL}(p^{*}||p^{t})\leq\frac{D_{KL}(p^{*}||p^{t-1})}{c}$. Applying this inequality recurrently we have $D_{KL}(p^{*}||p^{t})\leq\frac{D_{KL}(p^{*}||p^{0})}{c^{t}}\leq\frac{\log(m)}{c^{t}}$. Therefore, rearranging the terms in the (4) we have
$f(p^{t})+D_{KL}(q^{*}||q^{t})+D_{KL}(p^{*}||p^{t+1})=D_{KL}(p^{*}||p^{t}).$
Hence, $f(p^{t})\leq\frac{\log(m)}{c^{t}}$ which provides the desideratum.
Consequently, $f(p^{t})\to 0$ for the sequence $\{p^{t}\}$, hence the sequence converges to an optimal solution within $S$ with inverse exponential rate.
Finally, utilizing this inverse exponential bound, in a similar manner as in the proof of Theorem 4, we achieve the same bounds for the number of iterations and the total complexity, to achieve a solution [^5] Not necessary the solution $p^{*}$, since when the algorithm achieves an optimal solution stops. in $S$. ∎
The situation described in Theorem 5 is illustrated in Figure LABEL:fig:thm5.
5 Discussion
In this paper, we revisited the classical Arimoto-Blahut algorithm regarding the channel capacity problem and introduced significant analysis improvements that confirm its computational efficiency. Our new results give a mathematical guarantee to achieve an $\varepsilon$-optimal solution, for any constant $\varepsilon>0$, in substantially fewer iterations compared to the iterations required by the previous analyses. By scrutinizing the convergence analysis, we have established its reduced overall complexity, indicating that the Arimoto-Blahut algorithm can be indeed practical for large-scale applications. Proving advancements and theoretical insights about the performance of the algorithm, benefits problems where the Arimoto-Blahut algorithm is applied, such as the rate distortion problem or the Expectation-Maximization algorithm.
Acknowledgments
For this work the first and the second author were partially supported by the LASCON project of ICS-FORTH’s internal grants 2022.
[…]